A converging beam of rays is incident on a diverging lens.Having passed through the lens the rays intersect at a point 15cm from the lens on the opposite side.If the lens is removed the point where the rays will meet will move 5cm closer to the lens.The focal length of the lens is ?

Grade:10

2 Answers

Arun
6 years ago
Dear student
 
Given,
Image distance, v=15 cm
Object distance, u=15-5=10 cm (It is the distance when lens is removed. i.e object distance)
Now,
Using formula,
1/v-1/u=1/f
or 1/f=1/15 -1/10
or f=-30 cm
 
 
Regards
Arun (askIITians forum expert)
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Given,
Image distance, v=15 cm
Object distance, u=15-5=10 cm
(It is the distance when lens is removed. i.e object distance)
Now, Using formula,
1/v-1/u=1/f
or
1/f=1/15 -1/10
or
f=-30 cm

Thanks and Regards